OpenStudy (anonymous):

x^3 times x^-6 The answer must have a positive exponent I'm thinking x^18? THANK FOR YOU ANY HELP RECEIVED

OpenStudy (anonymous):

Does this really say x^3 times x^-6? I don't see a way to combine those.

OpenStudy (anonymous):

never mind 1/x^3

OpenStudy (anonymous):

x^a * x^b = x^(a+b) so x^3 times x^-6 = x^(3-6) = x^-3 = 1/x^3